OpenStudy (anonymous):

A line passes through the point (0, 3) and has a slope of 5/4 What is the equation of the line? A. 5x – 4y = –12 B. 5x + 4y = 12 C. 5x – 4y = –3 D. 5x + 4y = 3

OpenStudy (mathmale):

Please plug the given information into the "point-slope form of the equation of a straight line." Know that equation?

OpenStudy (anonymous):

use this equation (y-3) = m (x - 0)

OpenStudy (mathmale):

Yes, and that equation is a specific case of \[y-y _{0}=m(x-x _{0})\]...which is imporant and well worth knowing.